Circular No. 7225

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA AQUILAE 1999 T. Tomov, D. Moro, and U. Munari, Padova and Asiago Astronomical

Observatories, write: “Spectra (range 390–680 nm; 0.2 nm/pixel) of N Aql 1999 were obtained on July 15.86–15.94 UT with the Asiago 1.22-m telescope (+ Boller & Chivens CCD spectrograph). The FWHM of Hα, Hβ, Hγ, Hδ, and Hϵ emission lines are 3280, 3320, 3390, 3430, and 3440 km/s, and their equivalent widths are 80, 27, 8, 7, and 6 nm, respectively. Strong Fe II emission lines come from multiplet 42 at 492, 502, and 517 nm and multiplet 74 at 615, 624, and 645 nm. Other Fe II multiplets (37, 38, 48, 49, and 55) appear in emission, too. No line shows a convincing P-Cyg profile. Apart from a possible [O I] 558 nm, no strong forbidden line has yet appeared. The 0.25-nm equivalent width of the interstellar Na I doublet at 589 nm suggests a large reddening.”

Visual magnitude estimates: July 14.886 UT, 10.1 (K. Hornoch, Lelekovice, Czech Republic); 15.12, 10.4 (A. Pereira, Cabo da Roca, Portugal); 15.924, 10.9 (R. J. Bouma, Groningen, The Netherlands); 15.967, 10.9 (W. C. Souza, Joaquim Egidio, Brazil).

GRB 990712 G. Bakos, Space Telescope Science Institute (STScI) and Konkoly Ob-

servatory; K. Sahu, STScI; and J. Menzies, South African Astronomical Observatory (SAAO), on behalf of the PLANET collaboration; and P. Vreeswijk, University of Amsterdam, and F. Frontera, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna, on behalf of the Amsterdam and BeppoSAX/Bologna GRB follow-up teams, report: “The GRB 990712 WFC error circle (IAUC 7721) was observed with the 1-m SAAO telescope at Sutherland on July 12, 13, and 14. An R image taken on July 12.82 UT (integration time 900 s), when compared with the Digitized Sky Survey (DSS) image, showed a source at α = 22^{h}31^{m}53^{s}.1, δ = −73^{o}24′29′′ (equinox 2000.0; positional uncertainty 1′′) that is absent in the DSS image. Images taken during the following two nights showed the source to fade, and this further supports the identification. For photometric calibration, the standard stars F203 and F209 were observed, and the resultant magnitudes for the object are: July 12.82, R = 19.4; 13.74, R = 20.9; 13.76, I = 20.5; 14.76, R = 21.2; 14.77, V = 21.9. A finding chart is available at

http://www.stsci.edu/~ksahu/grb990712.html.”

1999 July 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green